OpenStudy (anonymous):

Find the product of -a 2b 2c 2(a + b - c)

OpenStudy (michele_laino):

please can you rewrite the expression using the editor?

OpenStudy (michele_laino):

is the expression like this: \[\large \left( { - a} \right) \cdot \left( {2b} \right) \cdot \left( {2c} \right) \cdot 2\left( {a + b - c} \right)\]